OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 + 5x3 - 2x2 + 5x - 3; d(x) = x2 + 1

OpenStudy (anonymous):

f(x) = (x2 + 1)( x2 + 5x - 3) f(x) = (x2 + 1)( x2 + 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3

OpenStudy (anonymous):

@hartnn im confused on this answer

OpenStudy (anonymous):

@musa19965